OpenStudy (anonymous):

A company manufactures computer chips. When testing the chips after they come off the production line, the company knows that a chip will be defective 10% of the time. What is the probability that, out of 50 chips, at least one is defective?

OpenStudy (kropot72):

The probability that zero chips out of 50 are faulty is given by: \[P(0\ defective\ out\ of\ 50)=\left(\begin{matrix}50 \\ 0\end{matrix}\right)(0.1)^{0}(0.9)^{50}=(0.9)^{50}\] \[P(at\ least\ 1\ defective)=1-(0.9)^{50}=you\ can\ calculate\]